# Universal Neurons in GPT-2: Emergence, Persistence, and Functional Impact

### v1

@@ Abstract
We investigate the phenomenon of neuron universality in independently trained GPT-2 Small models, examining how these universal neurons—neurons with consistently correlated activations across models—emerge and evolve throughout training. By analyzing five GPT-2 models at three checkpoints (100k, 200k, 300k steps), we identify universal neurons through pairwise correlation analysis of activations over a dataset of 5 million tokens. Ablation experiments reveal significant functional impacts of universal neurons on model predictions, measured via loss and KL divergence. Additionally, we quantify neuron persistence, demonstrating high stability of universal neurons across training checkpoints, particularly in deeper layers. These findings suggest stable and universal representational structures emerge during neural network training.
@@ 1 Introduction
Large language models (LLMs) exhibit remarkable generalization but remain difficult to interpret (Bommasani et al., 2021). However, neural networks are fully observable and deterministic, allowing us to record and manipulate internal components such as neuron activations (Bricken et al., 2023). This presents a rare opportunity to reverse-engineer their internal mechanism. An important open question regarding interpretability is whether models independently trained on the same task converge on similar internal structures—a notion termed the universality hypothesis (Wang et al., 2024). Universality, if established, offers stable interpretability targets and aids transfer learning.
We examine this hypothesis by analyzing five GPT-2 models trained from scratch, tracking when universal neurons—units with highly correlated activations across models (Gurnee et al., 2024)—emerge, their stability over training, and their causal role.
@@
  • Emergence Analysis: We provide the first systematic study of how universal neurons emerge during training, showing that they appear early and grow steadily, especially in deeper layers.
  • Persistence Quantification: We quantify the stability of universal neurons across training checkpoints, finding that over 80% remain universal in subsequent stages.
  • Functional Role via Ablation:We demonstrate that ablating universal neurons significantly increases loss and KL divergence, confirming their causal importance to model predictions.
@@ 2 Related Work
Early studies reported limited direct neuron matching (Kornblith et al., 2019). However, recent work identifies universal neurons with consistent semantic features across independently trained GPT-2 models (Gurnee et al., 2024). These universal neurons are also shown to correspond to interpretable, semantically meaningful features (Gurnee et al., 2024). This provides evidence that some circuits are consistently discovered across training runs, supporting the hypothesis of shared representational scaffolding.
@@
Because neurons are often polysemantic (Bricken et al., 2023), direct comparison is difficult. Lan et al. (2025) addressed this by learning sparse features with autoencoders and found significant alignment of feature dimensions across models. Algorithmic behaviors also show cross-architecture consistency, indicating broader universality (Lindsey et al., 2025; Ameisen et al., 2025).
@@
Works such as the lottery ticket hypothesis (Frankle & Carbin, 2019) and canonical correlation studies (Raghu et al., 2017) show that networks form persistent representational patterns within the first few training epochs. Theoretical analyses further support the idea that networks rapidly learn dominant features that are gradually refined (Saxe et al., 2019). These results motivate our focus on the emergence and persistence of universal neurons throughout training.
@@ 3 Method
We analyze five GPT-2 Small models at checkpoints 25%, 50%, and 75% of training(100k, 200k, 300k steps). Neuron activations are extracted over 5M tokens from the Pile dataset (monology, 2023).
@@
Following Gurnee et al. (2024), we compute Pearson correlations between neurons across model pairs. Let $\mathbf{a}_{k}^{(m,c)}\in\mathbb{R}^{n}$ denote the activation vector of neuron $k$ in model $m$ at checkpoint $c$ over $n$ token positions. The Pearson correlation between neurons defined by $\mathbf{a}_{k}^{(m_{1},c)}$ and $\mathbf{a}_{\ell}^{(m_{2},c)}$ is:
@@
where $\mu_{k}$ and $\sigma_{k}$ are mean and standard deviations of the activation vector $\mathbf{a}_{k}^{(m,c)}$ computed across a 5 million token dataset of the uncopyrighted Pile HuggingFace dataset (monology, 2023). We compute the excess correlation for a neuron $k$ with respect to a model $m_{2}$ at checkpoint $c$ as:
@@
where $\bar{\rho}^{m_{1},m_{2},c}_{k,\ell}$ is the pearson correlation between neuron $k$ in model $m_{1}$ and neuron $\ell$ in a randomly rotated version of the layer from model $m_{2}$, all at a checkpoint $c$. This rotation is constructed by multiplying the matrix of activations in that layer with a random Gaussian matrix, as described in Gurnee et al. (2024). The purpose of this transformation of activation vectors is to eliminate any privileged basis and establish a baseline for comparison (Gurnee et al., 2024). Neurons exceeding an excess correlation of 0.5 for a model $j$ are labeled universal. We also adjust this threshold to 0.4 and 0.6 to verify robustness. Tracking $\varrho_{i}$ across checkpoints and models allows us to observe universality emerge over training.
We used five GPT-2 Small models, models a through e 111from stanford-crfm at HuggingFace. We selected model a as the reference and computed Pearson correlations between its neurons and those in each of the other models (b, c, d, e). This yields four distinct sets of universal neurons.
@@
across training step intervals (e.g., 100$\rightarrow$200k, 200$\rightarrow$300k). To localize this further, we stratify by transformer layer $\ell$:
@@
To test functional significance, we ablate universal and control (non-universal) neurons during inference by zeroing their MLP outputs. We then measure changes in loss and KL divergence of the softmax distributions before and after ablation. We compute KL divergence as
$\mathrm{KL}(P||Q)=\frac{1}{|T|}\sum_{t\in T}\sum_{x\in V}P_{t}(x)\text{log}\frac{P_{t}(x)}{Q_{t}(x)}$
where $T$ is the set of token positions, $V$ is the output vocabulary (set of all possible tokens), and $P,Q$ are original and ablated softmax distributions of output logits at position $t$, respectively. We also perform a sensitivity analysis by repeating the experiment with relaxed thresholds for determining universality (e.g., 0.4 and 0.6).
@@ 4 Results
Universal neurons emerge early, increasing consistently through training, notably in deeper layers (Table 1). At early checkpoints (100k steps), fewer than 5% of neurons meet the universality criterion (0.5 threshold), but this fraction increases steadily to nearly 6% by 300k steps. Adjusting the universality threshold to 0.4 or 0.6 shows consistent trends.
Threshold 100k 200k 300k
0.4 9.58 10.99 11.33
0.5 4.74 5.56 5.71
0.6 2.00 2.35 2.41
Table 1: Percentage of universal neurons at different thresholds and checkpoints, averaged across models.
@@
We assess how consistently universal neurons remain universal as training progresses by computing their conditional persistence across three intervals: 100k→200k, 200k→300k, and 100k→300k steps.
Figure 2 shows layer-wise persistence rates across these intervals. We find that universal neurons are highly stable over time, especially in later layers. Layers 10 and 11 consistently exceed 90% persistence, while mid-layer neurons (e.g., layers 3–5) show greater volatility. Persistence from 100k→300k is slightly lower overall than adjacent intervals, reflecting gradual representational drift.
These results support the hypothesis that universal neurons, particularly in deeper layers, encode stable and task-relevant features that solidify as training proceeds. For completeness, we include detailed layer-wise persistence plots in Appendix A.1 that further proves the hypothesis.
@@
To test the functional importance of universal neurons, we ablate them by zeroing their MLP outputs during inference and measure the resulting change in model predictions using KL divergence.
Figure 2 shows that ablating all universal neurons (with excess correlation >0.5>0.5> 0.5) leads to a substantial shift in the output distribution, indicating a significant disruption in the model’s predictions. In contrast, ablating all non-universal neurons produces minimal change across all checkpoints and models.
These results demonstrate that universal neurons are not only shared across models but also causally important to inference. Considering that only about 5% of all neurons are universal, our results strongly support their role as core components of the model’s learned algorithm.
For completeness, we report supplementary ablation experiments in Appendix B. These include loss change and different thresholds (0.4 and 0.6). In addition to the above findings, layer-wise ablation(Appendix B.2) reveals that ablating universal neurons in the first layer causes a disproportionately large increase in both KL divergence and loss—far exceeding the impact observed in deeper layers. This suggests that early-layer universal neurons play a particularly critical role in shaping the model’s final predictions.
Figure 1: Persistence of universal neurons aggregated by layer (Y-axis begin from 60%).
Figure 2: KL divergence of ablated universal (solid) vs. non-universal (dashed) neurons. Colors indicate reference models.
@@ 5 Discussion and Conclusion
Findings In this paper, we explore how universal neurons - neurons with high correlations across models - emerge early on in training and persist throughout checkpoints. We found that universal neurons have high functional significance, as ablating them results in higher loss and KL divergence than non-universal neuron ablations. Neurons remain consistent across training checkpoints, with later layers having the higher persistence on average. Trends in universality continue to remain stable despite threshold adjustment (0.4 and 0.6).
@@
Future Work To further gain understanding of the impact of universal neurons across families, it would be interesting to examine ablations for families of universal neurons and how the loss varies. A wider selection of experiments could lead to greater insight, for instance, monitoring the effects of activation patching over some training data.
Appendix A Persistence of Universal Neurons over Checkpoints
A.1 Global Universal Neuron Persistence
Figure 3: Layer-aggregated persistence of universal neurons across training checkpoints for each model. Bars represent the percentage of universal neurons at the earlier checkpoint that remain universal at the later one. Later-stage intervals (e.g., 200k→300k) exhibit higher persistence, indicating stabilization of universal features over training.
@@ A.1 Global Universal Neuron Persistence
A.2 Detailed Layer-wise Universal Neuron Persistence
@@
All models exhibit a U-shaped trend: lower persistence in middle layers (2–5) and higher stability in both early and especially late layers. This suggests that early and late layers encode more stable, model-aligned features during training.
Figure 4: Layer-wise persistence of universal neurons from checkpoint 100k to 200k.
Figure 5: Layer-wise persistence of universal neurons from checkpoint 200k to 300k.
Figure 6: Layer-wise persistence of universal neurons from checkpoint 100k to 300k.
Appendix B Ablation Experiment Results
B.1 Loss Increase From Ablating All Universal/Non-universal Neurons
The increase in loss is computed as the difference in model loss before and after neuron ablation. Across all training checkpoints, ablating universal neurons results in a substantially greater increase in loss compared to non-universal neurons. The accompanying bar graphs support the primary claim presented in Section 4.
Figure 7: Loss increase from ablating universal neurons over training steps.
Figure 8: Loss increase from ablating non-universal neurons over training steps.
@@ B.1 Loss Increase From Ablating All Universal/Non-universal Neurons
B.2 Comparative Layer-Wise Effects of Universal vs. Non-Universal Neuron Ablation
At each training checkpoint, we ablate all universal or non-universal neurons within a specific layer and evaluate the resulting change in model output using KL divergence and loss increase. Figures below present the average KL divergence and loss difference (ablated minus original) across five models for both universal and non-universal neuron ablations.
Ablating universal neurons consistently leads to a greater increase in KL divergence and loss compared to non-universal neurons, indicating their stronger causal role in shaping model predictions. Notably, the first layer shows the most pronounced sensitivity to ablation, suggesting that early-layer universal neurons encode particularly critical information. In contrast, non-universal neuron ablation results in minor and often negligible effects across layers and checkpoints.
Figure 9: Layer-wise KL divergence from ablating universal neurons.
Figure 10: Layer-wise KL divergence from ablating non-universal neurons.
Figure 11: Layer-wise loss increase from ablating universal neurons.
Figure 12: Layer-wise loss increase from ablating non-universal neurons.
Appendix C Robustness Analysis Under Varying Universality Thresholds(Excess Correlation of 0.4 and 0.6)
Although the absolute values vary with different universality thresholds (0.4 and 0.6), the overall trends remain consistent. These results support our primary claims regarding the functional importance of universal neurons.
C.1 All Universal/Non-universal Neurons Ablation with 0.4 Thresholding
Figure 13: KL divergence from ablating universal neurons over training steps.
Figure 14: KL divergence from ablating non-universal neurons over training steps.
Figure 15: Loss increase from ablating universal neurons over training steps.
Figure 16: Loss increase from ablating non-universal neurons over training steps.
C.2 All Universal/Non-universal Neurons Ablation with 0.6 Thresholding
Figure 17: KL divergence from ablating universal neurons over training steps.
Figure 18: KL divergence from ablating non-universal neurons over training steps.
Figure 19: Loss increase from ablating universal neurons over training steps.
Figure 20: Loss increase from ablating non-universal neurons over training steps.

### v2

# Mechanistic Interpretability Workshop

@@ Abstract
We investigate the phenomenon of neuron universality in independently trained GPT-2 Small models, examining these universal neurons—neurons with consistently correlated activations across models—emerge and evolve throughout training. By analyzing five GPT-2 models at five checkpoints, we identify universal neurons through pairwise correlation analysis of activations over a dataset of 5 million tokens. Ablation experiments reveal significant functional impacts of universal neurons on model predictions, measured via cross entropy loss. Additionally, we quantify neuron persistence, demonstrating high stability of universal neurons across training checkpoints, particularly in early and deeper layers. These findings suggest stable and universal representational structures emerge during language model training.
@@ 1 Introduction
Large language models (LLMs) exhibit remarkable generalization but remain difficult to interpret (Bommasani2021FoundationModels). However, neural networks are fully observable and deterministic, allowing us to record and manipulate internal components such as neuron activations (bricken2023monosemanticity). This presents a rare opportunity to reverse-engineer their internal mechanism. An important open question regarding interpretability is whether models independently trained on the same task converge on similar internal structures—a notion termed the universality hypothesis (wang2024towards). Universality, if established, offers stable interpretability targets and aids transfer learning.
We examine this hypothesis by analyzing five GPT-2 models trained from scratch, tracking when universal neurons—units with highly correlated activations across models (gurnee2024universal)—emerge, their stability over training, and their causal role.
@@
  • Emergence Analysis: We provide the first systematic study on the emergence of universal neurons during training, showing that they appear early and grow steadily, especially in early and deeper layers.
  • Persistence Quantification: We quantify the stability of universal neurons across training checkpoints, finding that most remain universal in subsequent stages.
  • Functional Role via Ablation:We demonstrate that ablating universal neurons significantly increases loss, confirming their causal importance to model predictions.
@@ 2 Related Work
Early studies reported limited direct neuron matching (pmlr-v97-kornblith19a). However, recent work identifies universal neurons with consistent semantic features across independently trained GPT-2 models (gurnee2024universal). These universal neurons are also shown to correspond to interpretable, semantically meaningful features (gurnee2024universal). This provides evidence that some circuits are consistently discovered across training runs, supporting the hypothesis of shared representational scaffolding.
@@
Because neurons are often polysemantic (bricken2023monosemanticity), direct comparison is difficult. (lan2025sparseautoencodersrevealuniversal) addressed this by learning sparse features with autoencoders and found significant alignment of feature dimensions across models. Algorithmic behaviors also show cross-architecture consistency, indicating broader universality (lindsey2025biology; ameisen2025circuit).
@@
Works such as the lottery ticket hypothesis (frankle2018the) and canonical correlation studies (NIPS2017_7188) show that networks form persistent representational patterns within the first few training epochs. Theoretical analyses further support the idea that networks rapidly learn dominant features that are gradually refined (saxe2019mathematical). These results motivate our focus on the emergence and persistence of universal neurons throughout training.
@@ 3 Method
We analyze five GPT-2 Small models at checkpoints 20%, 40%, 60%, 80%, and 100% of training(80, 160k, 240k, 320k, 400k steps). Neuron activations are extracted over 5M tokens from the Pile dataset (monology_pile_uncopyrighted).
@@
Following (gurnee2024universal), we compute Pearson correlations between neurons across model pairs. Let $\mathbf{a}_{k}^{(m,c)}\in\mathbb{R}^{n}$ denote the activation vector of neuron $k$ in model $m$ at checkpoint $c$ over $n$ token positions. The Pearson correlation between neurons defined by $\mathbf{a}_{k}^{(m_{1},c)}$ and $\mathbf{a}_{\ell}^{(m_{2},c)}$ is:
@@
where $\mu_{k}$ and $\sigma_{k}$ are mean and standard deviations of the activation vector $\mathbf{a}_{k}^{(m,c)}$ computed across a 5 million token dataset of the uncopyrighted Pile HuggingFace dataset (monology_pile_uncopyrighted). We compute the excess correlation for a neuron $k$ with respect to a model $m_{2}$ at checkpoint $c$ as:
@@
where $\bar{\rho}^{m_{1},m_{2},c}_{k,\ell}$ is the pearson correlation between neuron $k$ in model $m_{1}$ and neuron $\ell$ in a randomly rotated version of the layer from model $m_{2}$, all at a checkpoint $c$. This rotation is constructed by multiplying the matrix of activations in that layer with a random Gaussian matrix, as described in (gurnee2024universal). The purpose of this transformation of activation vectors is to eliminate any privileged basis and establish a baseline for comparison (gurnee2024universal).
We used five GPT-2 Small models, models a through e 111from stanford-crfm at HuggingFace. We selected model a as the reference and computed Pearson correlations between its neurons and those in each of the other models (b, c, d, e). A neuron in model a is labeled universal if it averages an excess correlation above 0.5 across all 4 model pairs. We also adjust this threshold to 0.4 and 0.6 to verify robustness. Tracking $\varrho_{i}$ across checkpoints and models allows us to observe universality emerge over training.
@@
[…]
@@
To test functional significance, we ablate universal and control (non-universal) neurons during inference by zeroing their MLP outputs. We then measure changes in loss and change in loss per neuron ablated. We also perform a sensitivity analysis by repeating the experiment with relaxed thresholds for determining universality (e.g., 0.4 and 0.6).
@@ 4 Results
Universal neurons emerge early, increasing consistently through training, notably in earlier layers (Figure 1). At early checkpoints (80k steps), fewer than 5% of neurons meet the universality criterion (0.5 threshold), but this fraction increases steadily by 400k steps. As shown in Appendix A, adjusting the universality threshold to 0.4 or 0.6 shows consistent trends.
Figure 1: Percentage of Universal Neurons Across Layers. The graph shows an increasing trend of Universal Neurons as training step increases.
@@
We assess how universal neurons remain universal consistently as training progresses by computing their conditional persistence in five intervals: 80k→160k, 160k→240k, 240k→320k, 320k→400k, and 80k→400k steps.
Figure 2 shows layer-wise persistence rates across these intervals. We find that universal neurons are mostly highly stable over time, especially in early and later layers. Layers 0, 1, 10, and 11 maintain near or above 80% persistence, while mid-layer neurons (e.g., layers 3–8) show greater volatility. The overall persistence of 80k→400k is in general lower than the adjacent intervals, reflecting a gradual representational drift.
These results support the hypothesis that universal neurons, particularly in early and deeper layers, encode stable and task-relevant features that solidify as training proceed. For completeness, we include layer-wise persistence plots in Appendix B with different thresholds.
@@
To test the functional importance of universal neurons, we ablate them by zeroing their MLP outputs during inference and measure the resulting change in model predictions using Cross Entropy Loss. We measure the loss value of ablating groups of neurons and the ablation efficiency: change in loss per neuron ablated.
We perform four types of ablation experiments: ablating all universal neurons, ablating all non-universal neurons, ablating a random set of neurons equal in number to the universal neurons, and ablating five times as many random neurons. Figures 4&4 show that ablating all universal neurons (with excess correlation >0.5>0.5) leads to a substantial negative impact in the model’s predictions compared to ablating random neurons. Ablating universal neurons causes around the same level of disruption as ablating 5 times as many random neurons.
Figure 2: Universal Neuron Persistence Across Layers. Early and later layers show high Universal Neuron persistence, while middle layers experience shifting dynamics of universal features.
Figure 3: Absolute Loss Values After Ablating(zeroing activations) Different Neurons. It takes around 5x the amount of random neurons to achieve the same disruptive result of ablating universal neurons.
Figure 4: Ablation Efficiency (Change in loss per Neuron). The effect of universal neurons increases along with training steps. Compared to Nonuniversal neurons, they are crucial to the functionality of the language model.
Figure 5: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 80k.
Figure 6: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 400k.
These results demonstrate that universal neurons are not only shared across models but also causally important for inference. Considering that only about 4 to 5% of all neurons are universal, our results strongly support their role as core components of the model’s learned algorithm.
In addition to the above findings, layer-wise ablation(Figures 6&6) reveals that,throughout the training progress, ablating universal neurons in the first layer causes a disproportionately large increase in loss—far exceeding the impact observed in deeper layers. This suggests that early-layer universal neurons play a particularly critical role in shaping the model’s final predictions.
For completeness, we report comprehensive ablation experiments in Appendix C&D. These include global and layer-wise ablation results with different excess correlation thresholds (0.4 and 0.6).
@@ 5 Discussion and Conclusion
Findings In this paper, we explore how universal neurons - neurons with high correlations across models - emerge early on in training and persist throughout checkpoints. We found that universal neurons have high functional significance, as ablating them results in higher loss per neuron than non-universal neuron ablations. These universal neurons remain consistent across training checkpoints, with both early and later layers having higher persistence on average. Trends in universality continue to remain stable despite threshold adjustment (0.4 and 0.6).
@@
Future Work To better understand the impact of universal neurons across families, it would be interesting to examine ablations for families of universal neurons and how the loss varies. A wider selection of experiments could lead to greater insight, for example, monitoring the effects of activation patching on some training data.
@@
Appendix A Layer-wise Universal Neuron Percentage for Different Thresholds
Changing the threshold for labeling universal neuron shows similar trends.
Figure 7: Percentage of Universal Neurons(Excess correlation > 0.4) Across Layers throughout different checkpoints.
Figure 8: Percentage of Universal Neurons(Excess correlation > 0.6) Across Layers throughout different checkpoints.
Appendix B Persistence of Universal Neurons over Checkpoints for Different Thresholds
@@
All thresholds exhibit a U-shaped trend: lower persistence in middle layers (3–8) and higher stability in both early and especially late layers. This suggests that early and late layers encode more stable, model-aligned features during training.
Figure 9: Persistence of Universal Neurons (Excess correlation > 0.4) Across Layers throughout different checkpoints.
Figure 10: Persistence of Universal Neurons (Excess correlation > 0.6) Across Layers throughout different checkpoints.
Appendix C Additional Universal Neuron Ablation Experiment Results
As confirmed in Figure 12&14, universal neurons are more significant than nonuniversal neurons, but the magnitude differed greatly. The impact of ablating universal neurons when thresholding with 0.4 is greatly overwhelmed by ablating 5x the amount of random neurons. Whereas the impact of ablating universal neurons when thresholding with 0.6 is comparable to that of ablating the same number of random neurons.
However, Figure 12&14 consistently prove the functional significance of universal neurons through per-neuron metric.
C.1 Ablation Loss Increase with 0.4 Excess Correlation Thresholding
Figure 11: Absolute Loss Values After Ablating(zeroing activations) Different Neurons. When we lower the threshold from an excess correlation of 0.5 to that of 0.4, ablating 5x the amount of random neurons achieve more disruptive results than ablating all universal neurons.
Figure 12: Ablation Efficiency (Change in loss per Neuron). The effect of universal neurons increases along with training steps. Compared to Nonuniversal neurons, they are crucial to the functionality of the language model.
C.2 Ablation Loss Increase with 0.6 Excess Correlation Thresholding
Figure 13: Absolute Loss Values After Ablating(zeroing activations) Different Neurons. It takes around 5x the amount of random neurons to achieve the same disruptive result of ablating universal neurons.
Figure 14: Ablation Efficiency (Change in loss per Neuron). The effect of universal neurons increases along with training steps. Compared to Nonuniversal neurons, they are crucial to the functionality of the language model.
@@ Appendix C Additional Universal Neuron Ablation Experiment Results
Appendix D Additional Layer-Wise Universal Neuron Ablation Experiment Results
As shown in all figures below, the first layer of the model contains universal neurons that are the most crucial.
D.1 Layer-wise Ablation Efficiency Across Checkpoints
Figure 15: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 80k.
Figure 16: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 160k.
Figure 17: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 240k.
Figure 18: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 320k.
Figure 19: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 400k.
D.2 Layer-wise Ablation Efficiency with 0.4 Excess Correlation Thresholding
Figure 20: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 80k.
Figure 21: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 160k.
Figure 22: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 240k.
Figure 23: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 320k.
Figure 24: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 400k.
D.3 Layer-wise Ablation Efficiency with 0.6 Excess Correlation Thresholding
Figure 25: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 80k.
Figure 26: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 160k.
Figure 27: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 240k.
Figure 28: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 320k.
Figure 29: Layer-wise Ablation Efficiency (Change in loss per Neuron) on checkpoint 400k.